# Multi-Turn Code Generation Through Single-Step Rewards

### v1

@@ 4.3.4 Test-time scaling
Figure 4: Test-time scaling with different values of candidate solutions $N$ at each turn and different ways of learning verifiers. We compare with verifiers learned on samples from $\mu$Code and base policy. The candidate solutions are obtained from the 1B generator of $\mu$Code at each turn. We observe that the BoN performance improves with larger values of N on both datasets. The verifier learned with on-policy samples perform better.
@@
In the multi-turn setting, the number of candidate solutions can rise exponentially with the number of turns. To avoid this, $\mu$Code uses the learned verifier during inference to select the most promising candidate among $N$ candidates at each turn, leading to a linearly increasing number of calls to the generator. We study the inference-time scaling behaviors of $\mu$Code where we scale the number of candidate generations $N$ at each turn. Figure 4 plots the BoN with different values of $N$ ($1\leq N\leq 11$). The performance gains diminishes for larger $N$ on both datasets. On the MBPP dataset, the performance gains diminish with $N\geq 5$, whereas on HumanEval dataset a dip is observed for $N>7$.
In this section, we further study the importance of training the verifier with on-policy rollouts from the generator. We present a comparison of a verifier trained with samples from the Llama-3.2-1B-Instruct model (Base Verifier) and a verifier learned with samples from the generator of $\mu$Code. Note that we use the generator of $\mu$Code to obtain candidate solutions at each turn during evaluation. In Figure 4, we also present the scaling behaviors of different learned verifiers. We observe that using a verifier trained with on-policy samples obtained via the generator of $\mu$Code performs better and showed significant improvements of up to 4.3% for different values of candidate solutions $N$.
@@ 5 Related Work
Some work has explored explicitly training critics or reward models for multi-step reasoning tasks. In the coding domain, CodeRL (Le et al., 2022) trains a token-level critic to aid in code generation and to perform inference-time search. CodeRL’s mechanics are similar to our method, but their generator is not trained for multi-step: CodeRL trains a “code repairer” which conditions on one erroneous code completion while our generator incorporates multiple. ARCHER (Zhou et al., 2024), which frames multi-step tasks via a two-level hierarchical MDP, where the higher level MDP considers completions as actions and the lower level MDP considers tokens as actions. Another line of work utilizes Monte Carlo Tree Search (MCTS) methods for training: rStar-Math (Guan et al., 2025) trains a policy preference model to boost small LMs’ math abilities to match or exceed large reasoning-based LMs and ReST-MCTS (Zhang et al., 2024) trains a process reward model (PRM) similarly to Math-Shepherd (Wang et al., 2024a). Although $\mu$Code’s BoN search resembles a tree search, our key insight that multi-step code generation resembles a one-step recoverable MDP allows us to collect training trajectories much more efficiently. Finally, some work has explored using verifiers only during inference time. In “Let’s Verify Step by Step” (Lightman et al., 2023), the authors demonstrate that PRMs trained on erroneous math solutions annotated by humans outperform outcome reward models for filtering multiple inference time generations. Meanwhile, AlphaCode (Li et al., 2022) trains a test generator to evaluate multiple code solutions.
Other works omit learning a critic or reward model altogether. In the coding domain, RLEF (Gehring et al., 2024b) derives rewards only on the executor’s result on test cases and syntax checkers, and PPOCoder (Shojaee et al., 2023) additionally considers semantic and syntactic alignment, generated via data flow graphs and abstract syntax trees respectively, with a reference solution. The “oracle” rewards in these methods may not be informative for training, and in the case of PPOCoder, require complex constructs. We empirically show that having a reward model is beneficial by comparing $\mu$Code against the Multi-STaR baseline. Meanwhile, SCoRe (Kumar et al., 2024a) splits training into a “generator” and “correction” phase, thus restricting the total number of turns to 2. RISE (Qu et al., 2024) generates recovery steps via a more powerful LLM or by selecting a sampled completion via the oracle rewards. Both methods are less efficient than $\mu$Code, which doesn’t require generating corrections beyond generating training trajectories. Finally, FireAct (Chen et al., 2023a) and LEAP (Choudhury & Sodhi, 2024) FT ReAct style agents while RL4VLM (Zhai et al., 2024) and GLAM (Carta et al., 2024) studies training LLMs with interactive environment feedback.
@@ Acknowledgements
AJ is supported by Fonds de Recherche du Québec (FRQ), Calcul Québec, Canada CIFAR AI Chair program, and Canada Excellence Research Chairs (CERC) program. The authors are also grateful to Mila (mila.quebec) IDT and Digital Research Alliance of Canada for computing resources. AMR is supported in part by NSF CAREER #2037519 and NSF #2242302. SC is supported in part by Google Faculty Research Award, OpenAI SuperAlignment Grant, ONR Young Investigator Award, NSF RI #2312956, and NSF FRR#2327973.
Appendix A Appendix
@@
According to the one-step recoverable MDP definition, $A^{\star}(s,a)\leq 1$ for all $(s,a)$. Hence we can bound the performance difference as
@@
A.2 Hyperparameters
@@ A.2.2 Inference Parameters
A.3 Prompts
A.3.1 Single Step Prompt
Immediately below is the prompt template to generate 1 code completion in a single-step method or to generate the 1st step in a multi-step method. Below the prompt templates are examples of the code prompt and public tests for HumanEval and MBPP.
Single Step Prompt HumanEval Prompt HumanEval Test MBPP Prompt MBPP Test
A.3.2 Feedback Prompt
@@
A.4 Public Private Tests

### v2

@@ 4.3.4 Does $\mu$Code scale with inference budget?
Figure 3: Test-time scaling with different values of candidate solutions $N$ at each turn. The candidate solutions are obtained from the 1B-sized generator of $\mu$Code. We observe that the BoN performance improves with larger values of N on both datasets.
@@
In the multi-turn setting, the number of candidate solutions can rise exponentially with the number of turns. To avoid this, $\mu$Code uses the learned verifier during inference to select the most promising candidate among $N$ candidates at each turn, leading to a linearly increasing number of calls to the generator. In this experiment, we study the inference-time scaling behaviors of $\mu$Code where we scale the number of candidate generations $N$ at each turn. Figure 3 plots the BoN with different values of $N$ ($1\leq N\leq 11$). With more inference time budget, we observe that the performance improves with larger number of candidates at each turn on both datasets. The BoN accuracy plateaus with $N\geq 5$ for HumanEval dataset where for MBPP dataset we still observe some performance gains with larger $N$.
@@ 5 Related Work
Some work has explored explicitly training critics or reward models for multi-step reasoning tasks. In the coding domain, CodeRL (Le et al., 2022) trains a token-level critic to aid in code generation and to perform inference-time search. CodeRL’s mechanics are similar to our method, but their generator is not trained for multi-step: CodeRL trains a “code repairer” which conditions on one erroneous code completion while our generator incorporates multiple. ARCHER (Zhou et al., 2024), which frames multi-step tasks via a two-level hierarchical MDP, where the higher level MDP considers completions as actions and the lower level MDP considers tokens as actions. Another line of work utilizes Monte Carlo Tree Search (MCTS) methods for training: rStar-Math (Guan et al., 2025) trains a policy preference model to boost small LMs’ math abilities to match or exceed large reasoning-based LMs and ReST-MCTS (Zhang et al., 2024) trains a process reward model (PRM) similarly to Math-Shepherd (Wang et al., 2024a). Although $\mu$Code’s BoN search resembles a tree search, our key insight that multi-step code generation resembles a one-step recoverable MDP allows us to collect training trajectories much more efficiently. Finally, some work has explored using verifiers only during inference time. In “Let’s Verify Step by Step” (Lightman et al., 2023), the authors demonstrate that PRMs trained on erroneous math solutions annotated by humans outperform outcome reward models for filtering multiple inference time generations. Meanwhile, AlphaCode (Li et al., 2022b) trains a test generator to evaluate multiple code solutions.
Other works omit learning a critic or reward model altogether. In the coding domain, RLEF (Gehring et al., 2024b) derives rewards only on the executor’s result on test cases and syntax checkers, and PPOCoder (Shojaee et al., 2023) additionally considers semantic and syntactic alignment, generated via data flow graphs and abstract syntax trees respectively, with a reference solution. The “oracle” rewards in these methods may not be informative for training, and in the case of PPOCoder, require complex constructs. We empirically show that having a reward model is beneficial by comparing $\mu$Code against the RFT baseline. Meanwhile, SCoRe (Kumar et al., 2024a) splits training into a “generator” and “correction” phase, thus restricting the total number of turns to 2. RISE (Qu et al., 2024) generates recovery steps via a more powerful LLM or by selecting a sampled completion via the oracle rewards. Both methods are less efficient than $\mu$Code, which doesn’t require generating corrections beyond generating training trajectories. Finally, FireAct (Chen et al., 2023a) and LEAP (Choudhury & Sodhi, 2024) FT ReAct style agents while RL4VLM (Zhai et al., 2024) and GLAM (Carta et al., 2024) studies training LLMs with interactive environment feedback.
@@ Acknowledgements
AJ is supported by Calcul Québec, Canada Excellence Research Chairs (CERC), and Fonds de Recherche du Québec (FRQ) scholarship (DOI assigned: https://doi.org/10.69777/350253) program. The authors are also grateful to Mila (mila.quebec) IDT and Digital Research Alliance of Canada for computing resources. AMR is supported in part by NSF CAREER #2037519 and NSF #2242302. SC is supported in part by Google Faculty Research Award, OpenAI SuperAlignment Grant, ONR Young Investigator Award, NSF RI #2312956, and NSF FRR#2327973.
Appendix A Proofs
@@
According to the one-step recoverable MDP definition, $-1\leq A^{\star}(s,a)\leq 0$ for all $(s,a)$. Hence we can bound the performance difference as
@@
Appendix B Additional Results
B.1 Qwen Model
Method N MBPP HE
Instruct 1 53.8 64.5
    +BoN 5 60.9 70.3
RFT 1 57.0 66.6
    +BoN 5 58.0 71.3
$\mu$Code 1 59.0 70.5
    +BoN 5 63.1 74.0
Table 4: Comparison of our method $\mu$Code with baselines across MBPP, HumanEval using the Qwen-2.5-1.5B-Instruct model.
B.2 Additional private tests
Method N MBPP+ HE+
Instruct 1 43.2 25.7
    +BoN 5 49.9 31.9
RFT 1 44.3 26.7
    +BoN 5 50.0 34.3
$\mu$Code 1 48.7 32.4
    +BoN 5 55.1 40.0
Table 5: Comparison of our method $\mu$Code with baselines across MBPP+, HumanEval+ using the Llama-3.2-1B-Instruct model. MBPP+ and HumanEval+ introduce 35x and 80x more tests than their original counterparts respectively. Note that MBPP+ has a higher score than MBPP because there are 30% fewer problems
Appendix C Hyperparameters
@@ C.0.2 Inference Parameters
C.1 Prompts
C.1.1 Single Step Prompt
Immediately below is the prompt template to generate 1 code completion in a single-step method or to generate the 1st step in a multi-step method. Below the prompt templates are examples of the code prompt and public tests for HumanEval, MBPP, and CodeContests.
Single Step Prompt HumanEval Prompt Example HumanEval Test Example MBPP Prompt Example MBPP Test Example CodeContests Prompt Example CodeContests Test Example
C.1.2 Feedback Prompt
@@
C.2 Public Private Tests